Supplemental Operations Report 29Nov2025

[title Supplemental Operations Report – November 29th]

Date: 11/29/2025
Name of person filing report: Anderson Wilder
Reason for Report: Routine
Non-Nominal Systems: Crew car. Power system battery, inverters. Hab outer shell. Spirit Rover

Power system:
1) 200 hr maintenance (including oil, oil filter, and air filter change) performed on 11/23 at 441 hrs on clock
2) Current generator hours – 562.

Propane Readings:
Station Tank: 72%
Director Tank: 49%
Intern Tank: 59%
Generator Tank: 40%
Propane (generator) scheduled for Tuesday 12/2

Water:
Hab Static Tank – 550 gallons (above safety level 130 gallons)
GreenHab – 200 gallons (above safety level 60 gallons)
Outpost tank – 450 gallons
Septic Tanks emptied: Yes on 10/20/2025

Rovers:
Sojourner rover used: Yes (training)
Hours: 221.9
Beginning Charge: 100 %
Ending Charge: _ %
Currently Charging: Yes

Spirit Rover has a contactor issue so will not run, diagnosis in progress

Cars:
Hab Car used and why, where: To Hanksville for supplies
Crew Car used and why, where: To Hanksville for supplies, Grand Junction for crew pickup/drop off
General notes and comments: Crew car driver’s side front ball joint is bad

Summary of Internet: Nominal.
EVA suits and radios:
Suits: Nominal.
Comms: Nominal.
T-Echo EVA-link: Nominal.

Campus:Nominal.
Summary of Hab Operations: Nominal.
Summary of GreenHab Operations: Nominal.
Summary of SciDome Operations: Nominal.
Summary of Observatories Operations: Nominal.
Summary of RAM Operations: Cleaned. Nominal.
Summary of Outpost Operations: Nominal.
Summary of Health and Safety Issues: Nominal.

Supplemental Operations Report 22Nov2025

[title Supplemental Operations Report – November 22nd]

Date: 11/22/2025
Name of person filing report: Anderson Wilder
Reason for Report: Routine
Non-Nominal Systems: Crew car. Power system battery, inverters. HAB outer shell,

Power system:
1) 200 hr maintenance (including oil, oil filter, and air filter change) next expected on 11/23.
2) Current generator hours – 425.

Propane Readings:
Station Tank: 80%
Director Tank: 61%
Intern Tank: 63%
Generator Tank: 78%

Water:
Hab Static Tank – 550 gallons (above safety level 130 gallons)
GreenHab – 200 gallons (above safety level 60 gallons)
Outpost tank – 450 gallons
Septic Tanks emptied: Yes on 10/20/2025

Rovers:
Sojourner rover used: Yes (training)
Hours: 221.7
Beginning Charge: 100 %
Ending Charge: _ %
Currently Charging: Yes

Cars:
Hab Car used and why, where: To Hanksville for supplies
Crew Car used and why, where: To Hanksville for supplies, Grand Junction for crew pick up/supplies.
General notes and comments: Crew car driver’s side front ball joint is bad

Summary of Internet: Nominal.
EVA suits and radios:
Suits: Nominal.
Comms: Nominal.
T-Echo EVA-link: Nominal.

Campus:Nominal.
Summary of Hab Operations: Nominal.
Summary of GreenHab Operations: Nominal.
Summary of SciDome Operations: Nominal.
Summary of Observatories Operations: Robotic observatory failing to open. Motors are turning on, but there’s no movement. Working with Peter on next steps.
Summary of RAM Operations: Cleaned. Nominal.
Summary of Outpost Operations: Nominal.
Summary of Health and Safety Issues: Nominal.

Astronomy Report – <date – November 22nd>

[category  astronomy-report]

Report title: Astronomy Report
Crew #: 303
Position: Astronomer
Reported prepared by: Vikram Kothari
Date: 22-11-2025
SOL: 0

Non-nominal systems: N/A

Notes on non-nominal systems: N/A

Summary of Astronomy Operations:
Arrival time was after dark, so I was not able to get into the Musk Observatory to complete the baseline audit. Planning to complete the audit tomorrow.
Powered up the Astronomy Laptop. An update was automatically applied. Everything is nominal. Laptop is currently charging.
Peter will be setting up my account for ROCS-16

Sol Summary – <date – November 22nd>

Report title: Sol Summary Report
Crew #: 303
Position: Executive Officer
Report prepared by: Sydney Phillips
Date: 22-11-2025
Sol: 0

Summary Title: The goal in sight was to arrive and receive training.

Mission Status: Nominal. All systems operational.

Sol Activity Summary:

We all left our homes around 4am. We’ve traveled most of the day, landing in Grand Junction International Airport around 12:30pm. We all ate lunch at Denny’s in Grand Junction before meeting the previous Analog team at the airport to get the van. We then drove 2 hours to Hanksville and made our way to Mars Desert Research Station. Anderson completed our training at the station and we will continue the rest of it tomorrow.

LOOK AHEAD PLAN:
Crew Physical Status: All in Good Health + Spirits – A bit tired

EVAs to Be Completed: 0

Reports to Be Filed: Sol Summary, Journalist Report, Crew Photos, Green Hab Report, Operations Report, EVA reports, EVA Request(s)
Support Requested: Rover training support requested tomorrow. Food & water restock requested from previous crew.

Mission Plan – <Date – November 22nd>

The mission plan for Crew-303 follows below, showing the crew roles, projects, and EVA schedule.

Crew Composition:
Commander: Michael Andrews
Executive Officer: Sydney Phillips
Health and Safety Officer: Eli Henry
Crew Engineer: Aditya Kothari
Crew Artist: Sabrina Thompson
Crew Journalist: Aidan Westerland
Crew Astronomer: Vikram Kothari

Mission Projects:
Flight suit demonstration: Artist Thompson and XO Phillips have a demonstration suit (made of FR material) for functionality and durability.
Day-in-the-LIfe Content: Artist Thompson will be collecting footage and interviews of team members for STEM outreach.
Rover assembly and operation: Engineer Kothari will be assembling a small remote controlled rover at the station and will be deploying it on an EVA.
Astronomy: Astronomer Kothari has a research plan to capture a light curve of Algol during its primary eclipse on 11/24.
Drone/Timelapse Photos: Journalist Westerland and HSO Henry will be using a drone during EVAs and a tripod camera at night to capture additional footage.
3D Scanning: CO Andrews will be collecting rock samples and creating digital twins of them so they can be collecting without disturbing the environment.

EVA High-Level Plan:
11/23: two EVA teams will both visit Marble Ritual and Robert’s Rock Garden for EVA familiarization and rover familiarization.
11/24: two EVA teams will visit either Green Mars View or Somerville Overlook to gain experience in long-distance comms and longer duration EVAs
11/25: two EVA teams will choose a final mission that complements finishing their objectives: Sea of Shells, Candor Chasma, or Hab Ridge.

Astronomy Report – <date – Nov 22nd>

[category  astronomy-report]

Crew 303 Astronomy Report 22-11-2025
Name of person filing report: VIkram Kothari
SOL: 0

Non-nominal systems: N/A

Notes on non-nominal systems: N/A

Summary of Astronomy Operations:
Arrival time was after dark, so I was not able to get into the Musk Observatory to complete the baseline audit. Planning to complete the audit tomorrow.
Powered up the Astronomy Laptop. An update was automatically applied. Everything is nominal. Laptop is currently charging.
Peter will be setting up my account for ROCS-16

Supplemental Operations Report 15Nov2025

[title Supplemental Operations Report – November 15th]

Date: 11/15/2025
Name of person filing report: Anderson Wilder
Reason for Report: Routine
Non-Nominal Systems: Crew car. Power system battery, inverters. HAB outer shell,

Power system:
1) 200 hr maintenance (including oil, oil filter, and air filter change) next expected on 11/22.
2) Current generator hours – 320.

Propane Readings:
Station Tank: 80%
Director Tank: 64%
Intern Tank: 69%
Generator Tank: 54%

Water:
Hab Static Tank – 550 gallons (above safety level 130 gallons)
GreenHab – 200 gallons (above safety level 60 gallons)
Outpost tank – 450 gallons
Septic Tanks emptied: Yes on 10/20/2025

Rovers:
Sojourner rover used: No
Hours: 221.7
Beginning Charge: 100 %
Ending Charge: _ %
Currently Charging: Yes

Cars:
Hab Car used and why, where: To Hanksville for supplies
Crew Car used and why, where: To Hanksville for supplies, Grand Junction/Moab for crew pick up/supplies.
General notes and comments: Crew car driver’s side front ball joint is bad

Summary of Internet: Nominal.
EVA suits and radios:
Suits: Nominal.
Comms: Nominal, headsets seem to be having some issues. 4 of them currently tested and working completely, 1 for sure seems faulty, 2 seem intermittent
T-Echo EVA-link: Nominal.

Campus:Nominal.
Summary of Hab Operations: Nominal.
Summary of GreenHab Operations: Nominal.
Summary of SciDome Operations: Nominal.
Summary of Observatories Operations: Robotic observatory failing to open. Motors are turning on, but there’s no movement. Working with Peter on next steps.
Summary of RAM Operations: Cleaned. Nominal.
Summary of Outpost Operations: Nominal.
Summary of Health and Safety Issues: Nominal.

GreenHab Report – <date – November 11th>

Report title: GreenHab Report
Crew # 322
Position: GreenHab Officer
Report prepared by: Leszek Orzechowski
Date: 11-11-2025
Sol: 2

Environmental control (fan & heater): GH Heater: 60 F, GH Fan max 80 F, GH Fan min 75
Average temperatures (last 24h): 80 F
Maximum temperature (last 24h): 120 F
Minimum temperature (last 24h): 60F
Hours of supplemental light: : Supplemental lights are turned off
Water usage for crops: 8 liters
Daily water usage for research and/or other purposes: none
Time(s) of watering for crops: 10 am 7:30 PM
Changes to crops: new micro greens were set up – I only used already opened micro greens seeds
Soil Moisture Level: From moist to wet after watering the plants
Microgreen Growth Experiment: SpeCES experiment starter with 2 black trays inside the cabinet and 2 outside. The whole setup is no the Greenhab.
Space Nutrition Task: Part 1 Nutrines Catalog is well underway. We only have 2h of Internet a day (also for personal use not only mission related uses) so it is very challenging to work on it when you need to tap into the Internet for values. We will make progress but the mission profile is what it is 🙂

Harvest: no harvest to report

Sol Summary – <date – November 11th>

Report title: Sol Summary Report
Crew #: 322
Position: Co-Commander B
Report prepared by: Alex Kroll
Date: 11-11-2025
Sol: 2

Summary Title: Sol Summary Report

Mission Status: Nominal

Sol Activity Summary:

2 EVAs
Mission Task Updates:

Space Nutrition Task: Working on task 1.2 – nutritional documentation
VR Training: 2 crewmembers complete
Mission Vlog: Initial planning completed, gathering footage
Ultrasound: All crewmembers trained, ready for individual trials
Look Ahead Plan:

Sol 3: 2 EVAs, self-exam US session for all AAs, continue VR training, space nutrition task.
Implement EVA cable management & time management plan, water conservation plan.

Anomalies in work:

General:
High Water usage – plan developed to reduce. Will take manual readings daily to verify.
Determining solution to cable management. Will brief solution when resolved.
Improve time management in EVAs, particularly 5min checkins
EVA 1:
NSTR pre-mission
Good comms on radio 2 & 7
Mission success
Tobii USB-C from powerbank connection broken upon return – prompts improvement in cable management
NSTR post-mission
Mission duration ~40min
EVA 2:
NSTR pre-mission – reduced equipment, dual-channel personal microphones on EV1, single channel on EV2.
Poor comms on radio 2 & 7
Tricorder 1 & 2 – No GPS signal for first 5-7 min
Digital compass unreliable
Mission presumed not successful
NSTR post-mission
Mission duration ~37min
Weather: Sunny with few clouds, maximum 64°F, minimum 30°F

Crew Physical Status: High spirited, established in roles, becoming more comfortable as a crew

EVA: 2

Reports to be filled: Sol Summary, Journalist Report, GreenHab Report, Operations Report, HSO Report

Support Requested: N/A

Sol Summary – December 22nd

Report title: Sol Summary Report
Crew #: 325
Position: Crew Commander
Report prepared by: Cesare Guariniello
Date: 22Dec2025
Sol: 7

Summary Title:
Mission Status: Nominal
Sol Activity Summary: With the slow rhythm of Mars, we woke up once again at 8am to have breakfast while having our morning communications. At 11, Commander Cesare, Crew Scientist Ellenah, and Crew Engineer Morgan suited up for an EVA in the northwestern region of MDRS, called “The Moons” in some old map. When the rovers reached the “do not proceed further” charge level, the EVA crew began setting up two experiments: Ellenah took readings from her solar power experiment at the highest elevation reached until now. Morgan began digging a trench for his experiment, supported by Cesare when possible. Ellenah and Cesare occupied the time when Morgan needed to work alone with some meditation and relax. After refilling the trench, the trio walked along Brahe Highway, then down to the Sea of Shells, where Cesare collected some gypsum samples. Back at the hab, the crew continued their work on projects. The afternoon brought a mix of work, tired naps, and a short movie. Working on our reports right now, while the Crew Journalist and GHO prepare dinner. We passed the halfway point of the mission, already on our trajectory back to Earth.
Look Ahead Plan: Weather should be good again tomorrow, good for a nice EVA. Getting ready for Holiday celebrations
Anomalies in work: None
Weather: Getting more cloudy
Crew Physical Status: Everybody is doing great
EVA: good exploration of the northwestern regions: higher elevation around north of Brahe, then down to the Sea of Shells
Reports to be filed: Sol Summary, Journalist Report, Greenhab Report, Photos (6-8 images), Operations Report, EVA 8 Report, EVA 9 Request.
Support Requested: None

Copyright © The Mars Society. All rights reserved. | Main Site